# Stochastic Conditional Diffusion Models for Robust Semantic Image Synthesis

### v2

@@ 7 Conclusion
In this paper, we introduce a novel and robust conditional diffusion model, SCDM, for semantic image synthesis. The discrete diffusion for labels, which we name Label Diffusion, is designed with label masking. Label Diffusion ensures that the intermediate labels along the generation process become similar and eventually identical at $t=T$. Additionally, the class-wise noise schedule improves the generation quality for small and rare objects. We define the generation process with a discrete forward process of labels and a continuous reverse process of images, as the labels are given in SIS. SCDM demonstrates its robustness in noisy SIS setups which we designed to simulate human errors in real-world applications.
Broader Impact
This paper focuses on robust generation against noisy user input, which is a crucial aspect to regard in real-world applications. As the clean labels and the benchmark dataset intrinsically contain biases, our model might reinforce the biases when the samples from our model are used prevalently throughout the Internet. Nevertheless, our research opens up new avenues for future works in image generation for real-world problems, potentially expanding the application of diffusion models in practical scenarios.
@@ A.3 Proof of Proposition 2
[…]
@@ Appendix B Algorithms
Algorithm 1 and 2 summarize the general training and sampling process of our SCD, respectively. While $\mathbf{m}$ has to be $\mathbf{e}_{C+1}\in\mathbb{R}^{C+1}$, we implemented the absorbing vector with a zero vector $\mathbf{0}\in\mathbb{R}^{C}$ to minimize the modification of the pretrained SIS model.
@@
0:  $\text{label }\mathbf{y}_{0}\in\mathbb{R}^{H\times W\times C},~{}\text{noise schedule }\alpha_{1:T},\gamma_{1:T},\newline \text{guidance scale }s,\text{absorbing vector }\mathbf{m}\in\mathbb{R}^{C},\newline \text{extrapolation scale }w$
@@
11:     $\mathbf{x}_{t-1}\leftarrow\mu_{t}+\Sigma_{\theta}(\mathbf{x}_{t},\mathbf{y}_{t},t)^{\frac{1}{2}}\mathbf{z}$// Sample $\mathbf{x}_{t-1}$}
@@ C.1 Metrics
Additionally, following previous works, we assess the alignment of the synthesis results with ground truth semantic maps by using off-the-shelf pretrained segmentation networks and report mean Intersection-over-Union (mIoU). We feed the sampled images to the pre-trained, off-the-shelf semantic segmentation networks: U-Net (Lee et al., 2020; Ronneberger et al., 2015) for CelebAMask-HQ, ViT-Adapter-S (Chen et al., 2022) with UperNet (Xiao et al., 2018) for ADE20K, and DeepLabV2 (Chen et al., 2015) for COCO-Stuff. For a fair comparison, we tried to measure and reproduce the mIoU of all baseline samples. For the standard SIS setting, when samples or checkpoints were not publicly available, we report the results from (Tang et al., 2023) and denote them with ‘*’ in the table.
C.1.2 Experiments in Section F.1
@@
In addition to FID and mIoU, we adopt Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as our generation diversity metrics in Section F.1 and compare the average distance between multi-modal synthesis results. Specifically, following (Tan et al., 2021a), we sample 10 different images for each semantic map, compute the pairwise LPIPS distance, and average over the label maps.
C.1.3 Experiments in Section 6.1.2
@@ C.2 Baselines
We compare our model with GAN-based methods such as pix2pixHD (Wang et al., 2018), SPADE (Park et al., 2019), DAGAN (Tang et al., 2020a), SCGAN (Wang et al., 2021), CLADE (Tan et al., 2021b), CC-FPSE (Liu et al., 2019), GroupDNet (Zhu et al., 2020b), INADE (Tan et al., 2021a), OASIS (Sushko et al., 2020), RESAIL (Shi et al., 2022), SAFM (Lv et al., 2022), and ECGAN (Tang et al., 2023). We also evaluate our method with diffusion-based approaches, SDM (Wang et al., 2022b), LDM (Rombach et al., 2022), and PITI (Wang et al., 2022a). Although FLIS (Xue et al., 2023) does tackle SIS with diffusion models, they utilize extra text inputs (captions) with the semantic maps. For a fair comparison, we did not include FLIS as a baseline.
As the pretrained weights of pix2pixHD and PITI on ADE20K are not publicly available, we excluded them from our noisy SIS experiments on ADE20K. Additionally, the pretrained weights and the full code of RESAIL and ECGAN are not available in public, therefore we did not include them in our noisy SIS experiments baselines as well.
@@ Appendix D Detailed Explanations on $\psi_{c}$ and $\phi_{c}$
In this section, we elaborate on $\psi_{c}$ and $\phi_{c}$ as introduced in Section 4.2 and provide some visual aid for $\gamma_{t,c}$. $\psi_{c}$ accounts for the area ($\approx$ object size) of class $c$. Eq. (15) defines $\psi_{c}$, which we estimated with training data. $\psi_{c}$ ensures that classes with smaller average areas transition to the masked state later. Specifically, the term is defined as the inverse of the ratio of average total image pixels to the average class pixels. As the average area of class $c$ decreases, the $\psi_{c}$ value diminishes. On the other hand, $\phi_{c}$ ensures a slow diffusion (to the masked state) of classes that rarely appear in the dataset. It is defined in Eq. (16), and it can be regarded as the inverse frequency of the class in training images. A larger $\phi_{c}$ value corresponds to a class that rarely appears in the dataset. In Figure 5, we visualize $\gamma_{t,c}$ with different noise schedules. The class-wise schedule of class ‘clock’ ensures that the class is diffused at a relatively later time, as its $\psi_{c}\phi_{c}$ value is large.
Appendix E ADE20K Dataset Annotation Examples
@@
Appendix F Additional Experimental Results
F.1 Standard SIS Setting
Table 5: Quantitative performance comparison on generation quality. The baseline methods are categorized into Generative Adversarial Networks (GAN) and Diffusion Models (DM). For FID, lower is better. For LPIPS and mIoU, higher is better. The best results among the diffusion-based approaches are boldfaced, and the best results overall are underlined. ‘-’ indicates that the method did not report the metric or did not train the dataset, or the checkpoint or samples are not publicly available. ‘Seg’ denotes that the method leverages a pre-trained segmentation network during training.
Methods Seg CelebAMask-HQ ADE20K COCO-Stuff
GAN RESAIL (Shi et al., 2022) - - - 30.2 - 49.3* 18.3 - 44.7
SAFM (Lv et al., 2022) - - - 32.8 - 52.6 24.6 - 43.3
ECGAN (Tang et al., 2023) - - - 25.8 0.52 50.6* 15.7 - 46.3
pix2pixHD (Wang et al., 2018) 38.5 0 76.1 81.8 0 20.3* 111.5 0 14.6
SPADE (Park et al., 2019) 29.2 0 75.2 33.9 0 44.5 33.9 0 36.9
CC-FPSE (Liu et al., 2019) - - - 31.7 0.078 47.3 19.2 0.098 40.8
DAGAN (Tang et al., 2020a) 29.1 0 76.6 31.9 0 45.5 - - -
GroupDNet (Zhu et al., 2020b) 25.9 0.365 76.1 41.7 0.230 33.7 - - -
OASIS (Sushko et al., 2020) - - - 28.3 0.286 50.9 17.0 0.328 44.2
INADE (Tan et al., 2021a) 21.5 0.415 74.1 35.2 0.459 41.4 - - -
SCGAN (Wang et al., 2021) 20.8 0 75.5 29.3 0 50.0 18.1 0 41.7
CLADE (Tan et al., 2021b) 30.6 0 75.4 35.4 0 44.7 29.2 0 36.9
LDM (Rombach et al., 2022) 21.5 0.315 74.6 36.5 0.417 23.2 - - -
PITI (Wang et al., 2022a) - - - 27.3 - - 15.8 0.489 32.2
Ours 17.4 0.418 77.2 26.9 0.530 49.4 15.3 0.519 37.9
F.1.1 Generation quality comparison.
We also evaluate our method in a standard SIS setting with clean labels. Table 5 shows that SCDM achieves the best performance in all three metrics (e.g., FID, LPIPS, mIoU) compared to recent DM-based baselines in all datasets. Also, including GAN-based models, the proposed method shows the best overall performance. Specifically, our method surpasses all baselines on CelebAMask-HQ in all three metrics with a significant gain of +1.4 (FID) compared to the state-of-the-art method.
@@ F.1 Standard SIS Setting
Figure 8: Qualitative results on COCO. Our SCD clearly captures the details (e.g., the hands on the clock) and naturally depicts the branches over the clock tower.
Figure 9: Robustness to inconsistent semantic labels. Our SCD exhibits robustness to inconsistent semantic labels. Although the parasol (class ‘umbrella’) in the original image has a shaft, it is not accurately annotated in the semantic map. Interestingly, ours synthesized a natural image with the missing shaft whereas all baselines failed to generate it.
F.1.2 Analysis on semantic correspondence.
Furthermore, we evaluate the semantic correspondence using pre-trained segmentation networks and compare our method with baselines on mIoU. SCDM shows better performance than the state-of-the-art on CelebAMask-HQ, while for the other two datasets, diffusion-based approaches including ours show comparatively weaker performance than adversarial methods such as SAFM or ECGAN. We believe that one of the reasons for this could be the erroneous annotation of images. For example, ‘umbrella’ is often annotated without a shaft and a handle (examples are given in Appendix E). In Figure 9, baselines trained with these images did not synthesize the shaft for the given semantic map of a parasol (class ‘umbrella’) without a shaft, while ours did. Although ours is more realistic, this can result in lower mIoU scores. We also suspect that the performances of the off-the-shelf segmentation models used for mIoU evaluation are not very robust. We ran the segmentation model on ground-truth (i.e., real and clean) images and it shows a poor mIoU with semantic labels (43.1 on ADE20K and 34.7 on COCO). Additionally, it is noteworthy that baselines with high mIoU scores such as RESAIL, SAFM, and ECGAN leverage ‘pretrained’ segmentation models in their semantic alignment losses (denoted by ‘Seg’ in Table 5). In contrast, ours does not rely on such explicit guidance for correspondence.
F.2 Extrapolation Hyperparameter Search
@@
F.3 Visualization of Label Diffusion
@@
F.4 Effect of Class-wise Noise Schedule
Table 7: Quantitative noise schedule comparison on semantic correspondence (mIoU (normal-$\uparrow$)). The results are sampled over 1000 sampling steps.
@@ F.4 Effect of Class-wise Noise Schedule
F.5 Further Discussion and Analysis on Class Guidance
@@
F.6 Validation of SCDM Generation Process
@@
F.7 More Qualitative Results - Multimodal Generation
@@
F.8 More Qualitative Results - SIS with Noisy Labels
@@
F.9 More Qualitative Results - Standard SIS Setting
@@
Figure 20: More qualitative results and comparisons on ADE20K. The first two row shows the result of our model, while the other rows depict qualitative comparisons in terms of generation quality.
Figure 21: More qualitative results and comparisons on COCO-Stuff. The first two row shows the result of our model, while the other rows depict qualitative comparisons in terms of generation quality.

### v3

@@ 7 Conclusion
This paper introduces SCDM, a novel and robust conditional diffusion model for semantic image synthesis. The discrete diffusion for labels, which we name Label Diffusion, is designed with label masking. Label Diffusion ensures that the intermediate labels along the generation process become similar and eventually identical at $t=T$. Additionally, the class-wise noise schedule improves the generation quality for small and rare objects. We define the generation process with a discrete forward process of labels and a continuous reverse process of images, as the labels are given in SIS. SCDM demonstrates its robustness in noisy SIS setups which we designed to simulate human errors in real-world applications.
Acknowledgements
This work was partly supported by ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT)(IITP-2024-2020-0-01819, 10%), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(NRF-2023R1A2C2005373, 30%), the Virtual Engineering Platform Project (Grant No. P0022336, 30%), funded by the Ministry of Trade, Industry & Energy (MoTIE, South Korea), and the National Supercomputing Center with supercomputing resources including technical support (KSC-2022-CRE-0261, 30%).
Impact Statement
This paper focuses on robust generation against noisy user input, a crucial aspect in real-world applications. Since clean labels and benchmark datasets inherently contain biases, our model might reinforce these biases when its samples are widely used across the Internet. Nevertheless, our research opens new avenues for future work in image generation for real-world problems, potentially expanding the application of diffusion models in practical scenarios.
@@ A.3 Proof of Proposition 2
Suppose there exists a differentiable function $f_{i}$ such that $q(\mathbf{y}_{0}^{i}|\mathbf{x}_{t})=\text{Cat}(\mathbf{y}_{0}^{i};p=f_{i}(\mathbf{x}_{t}))$ and $\mathbf{y}_{0}^{1}|\mathbf{x}_{t},...,\mathbf{y}_{0}^{H\times W}|\mathbf{x}_{t}$ are independent, where $i\in\{1,...,H\times W\}$ denotes the index of a pixel in semantic map and $\mathbf{y}_{0}^{i}\in\mathbb{R}^{C+1}$ is a one-hot vector. With $q(\mathbf{y}_{t}^{i}|\mathbf{y}_{0}^{i})$ from Eq. (13) and $\gamma_{t,c}=\gamma_{t}\text{ for any }c$, we have the following relationship;
@@ Appendix B Algorithms
Algorithm 1 and 2 summarize the general training and sampling process of our SCDM, respectively. While $\mathbf{m}$ has to be $\mathbf{e}_{C+1}\in\mathbb{R}^{C+1}$, we implemented the absorbing vector with a zero vector $\mathbf{0}\in\mathbb{R}^{C}$ to minimize the modification of the pretrained SIS model.
@@
0:  $\text{label }\mathbf{y}_{0}\in\mathbb{R}^{H\times W\times C},~{}\text{noise schedule }\alpha_{1:T},\gamma_{1:T},\text{guidance scale }s,\text{absorbing vector }\mathbf{m}\in\mathbb{R}^{C},\newline \text{extrapolation scale }w$
@@
11:     $\mathbf{x}_{t-1}\leftarrow\mu_{t}+\Sigma_{\theta}(\mathbf{x}_{t},\mathbf{y}_{t},t)^{\frac{1}{2}}\epsilon_{\mathbf{x}}$// Sample $\mathbf{x}_{t-1}$}
@@ C.1 Metrics
Additionally, following previous works, we assess the alignment of the synthesis results with ground truth semantic maps by using off-the-shelf pretrained segmentation networks and report mean Intersection-over-Union (mIoU). We feed the sampled images to the pre-trained, off-the-shelf semantic segmentation networks: U-Net (Lee et al., 2020; Ronneberger et al., 2015) for CelebAMask-HQ, ViT-Adapter-S (Chen et al., 2022) with UperNet (Xiao et al., 2018) for ADE20K, and DeepLabV2 (Chen et al., 2015) for COCO-Stuff. For a fair comparison, we tried to measure and reproduce the mIoU of all baseline samples. For the standard SIS setting, when samples or checkpoints were not publicly available, we report the results from (Tang et al., 2023a) and denote them with ‘*’ in the table.
@@
In addition to FID and mIoU, we adopt Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) as our generation diversity metrics in Section G.1 and compare the average distance between multi-modal synthesis results. Specifically, following (Tan et al., 2021a), we sample 10 different images for each semantic map, compute the pairwise LPIPS distance, and average over the label maps.
C.1.2 Experiments in Section 6.1.2
@@ C.2 Baselines
We compare our model with GAN-based methods such as pix2pixHD (Wang et al., 2018), SPADE (Park et al., 2019), DAGAN (Tang et al., 2020a), SCGAN (Wang et al., 2021), CLADE (Tan et al., 2021b), CC-FPSE (Liu et al., 2019), GroupDNet (Zhu et al., 2020b), INADE (Tan et al., 2021a), OASIS (Sushko et al., 2020), RESAIL (Shi et al., 2022), SAFM (Lv et al., 2022), ECGAN (Tang et al., 2023a), ECGAN++ (Tang et al., 2023b), and Unconditional Generator with Semantic Mapper (Chae et al., 2024). We also evaluate our method with diffusion-based approaches, SDM (Wang et al., 2022b), LDM (Rombach et al., 2022), and PITI (Wang et al., 2022a). Although FLIS (Xue et al., 2023) and ControlNet (Zhang et al., 2023) do tackle SIS with diffusion models, they utilize extra text inputs (captions) with the semantic maps. For a fair comparison, we did not include FLIS and ControlNet as baselines.
As the pretrained weights of pix2pixHD and PITI on ADE20K are not publicly available, we excluded them from our noisy SIS experiments on ADE20K. Additionally, the pretrained weights and the full code of RESAIL, ECGAN, and Unconditional Generator are not available in public, therefore we did not include them in our noisy SIS experiments baselines as well.
@@ Appendix D Detailed Explanations on $\psi_{c}$ and $\phi_{c}$
In this section, we elaborate on $\psi_{c}$ and $\phi_{c}$ as introduced in Section 4.2 and provide some visual aid for $\gamma_{t,c}$. $\psi_{c}$ accounts for the area ($\approx$ object size) of class $c$. Eq. (15) defines $\psi_{c}$, which we estimated with training data. $\psi_{c}$ ensures that classes with smaller average areas transition to the masked state later. Specifically, the term is defined as the inverse of the ratio of average total image pixels to the average class pixels. As the average area of class $c$ decreases, the $\psi_{c}$ value increases. On the other hand, $\phi_{c}$ ensures a slow diffusion (to the masked state) of classes that rarely appear in the dataset. It is defined in Eq. (16), and it can be regarded as the inverse frequency of the class in training images. A larger $\phi_{c}$ value corresponds to a class that rarely appears in the dataset. In Figure 5, we visualize $\gamma_{t,c}$ with different noise schedules. The class-wise schedule of class ‘clock’ ensures that the class is diffused at a relatively later time, as its $\psi_{c}\phi_{c}$ value is large.
Appendix E Limitations
One possible limitation of this work is that as the method aims to generate clean images from the noisy labels, the user’s intention might be ignored, i.e., the user might actually want to generate noisy images. Controlling faithfulness to the semantic maps could be one possible future direction, which we leave as future work. Another limitation could be that our method cannot dynamically learn the optimal noise schedules for labels. If the proposed method can learn the optimal noise schedule, it will improve the flexibility and power of semantic image synthesis.
Appendix F ADE20K Dataset Annotation Examples
@@
Appendix G Additional Experimental Results
G.1 Standard SIS Setting
G.1.1 Generation quality comparison.
@@ G.1 Standard SIS Setting
Figure 8: Qualitative results on COCO. Our SCDM clearly captures the details (e.g., the hands on the clock) and naturally depicts the branches over the clock tower.
Figure 9: Robustness to inconsistent semantic labels. Our SCDM exhibits robustness to inconsistent semantic labels. Although the parasol (class ‘umbrella’) in the original image has a shaft, it is not accurately annotated in the semantic map. Interestingly, ours synthesized a natural image with the missing shaft whereas all baselines failed to generate it.
G.1.2 Analysis on semantic correspondence.
We evaluate the semantic correspondence using pre-trained segmentation networks and compare our method with baselines on mIoU, and the results are in Table 2. SCDM shows better performance than the state-of-the-art on CelebAMask-HQ, while for the other two datasets, diffusion-based approaches including ours show comparatively weaker performance than adversarial methods such as SAFM or ECGAN. We believe that one of the reasons for this could be the erroneous annotation of images. For example, ‘umbrella’ is often annotated without a shaft and a handle (examples are given in Appendix F). In Figure 9, baselines trained with these images did not synthesize the shaft for the given semantic map of a parasol (class ‘umbrella’) without a shaft, while ours did. Although ours is more realistic, this can result in lower mIoU scores. We also suspect that the performance of the off-the-shelf segmentation models used for mIoU evaluation is not very robust. We ran the segmentation model on ground-truth (i.e., real and clean) images and it shows a poor mIoU with semantic labels (43.1 on ADE20K and 34.7 on COCO). Additionally, it is noteworthy that baselines with high mIoU scores such as RESAIL, SAFM, and ECGAN leverage ‘pretrained’ segmentation models in their semantic alignment losses (denoted by ‘Seg’ in Table 2). In contrast, ours does not rely on such explicit guidance for correspondence.
G.2 Extrapolation Hyperparameter Search
@@
G.3 Visualization of Label Diffusion
@@
G.4 Effect of Class-wise Noise Schedule
Table 7: Quantitative noise schedule comparison on semantic correspondence (mIoU ($\uparrow$)). The results are sampled over 1000 sampling steps.
@@ G.4 Effect of Class-wise Noise Schedule
G.5 Further Discussion and Analysis on Class Guidance
@@
G.6 Validation of SCDM Generation Process
@@
G.7 More Qualitative Results - Multimodal Generation
@@
G.8 More Qualitative Results - SIS with Noisy Labels
@@
G.9 More Qualitative Results - Standard SIS Setting
@@
Figure 20: More qualitative results and comparisons on ADE20K. The first two rows show the result of our model, while the other rows depict qualitative comparisons in terms of generation quality.
Figure 21: More qualitative results and comparisons on COCO-Stuff. The first two rows show the result of our model, while the other rows depict qualitative comparisons in terms of generation quality.